Demand for Kristal blanket is given by D_P = 42 − 2p , the supply function is

given by S_P = 5p;


What is the equilibrium price, output, consumer and producer surplus in the market for Kristal blankets?


1
Expert's answer
2021-11-17T10:01:10-0500

Solution:

At equilibrium: Qd = Q"\times"

42 – 2p = 5p

42 = 5p + 2p

42 = 7p

P = 6

Equilibrium price = 6

Substitute in either the demand or supply function to determine equilibrium quantity:

Qd = 42 – 2p = 42 – 2(6) = 42 – 12 = 30

Qs = 5p = 5 "\times" 6 = 30

Equilibrium quantity = 30

 

Consumer surplus = ½ "\times" (30)( 21 – 6) = 15 "\times" 15 = 225

 

Producer surplus = ½ "\times" (30) (6-2) = 15 "\times" 4 = 60